OpenStudy (anonymous):

the equation for the horizontal distnace h in feet of a projectile with initial velocity v(subscript o) and initial angle θ is given by h=v(subscrit 0)^2/16sinθcosθ Assume the initial velocity is 60 feet/seconds. What initial angle will you need to ensure that the horizontal distnace will be exactly 100 feet?

OpenStudy (dumbcow):

\[v_{o} = 60, h = 100\] \[100 = \frac{60^{2}}{16\sin(\theta)\cos(\theta)}\] Also use trig identity \[\sin(2\theta) = 2\sin(\theta)\cos(\theta)\] Substitute this in \[\rightarrow 100 = \frac{60^{2}}{8\sin(2\theta)}\] \[800\sin(2\theta) = 60^{2}\] \[\sin(2\theta) = \frac{60^{2}}{800} = 4.5\] hmm that cant be right, range of sin is -1 to 1 did you copy the equation correctly ??

OpenStudy (dumbcow):

maybe its \[h = \frac{v_{o}\sin(\theta)\cos(\theta)}{16}\] ??

OpenStudy (dumbcow):

oh well, i'll assume its \[h = \frac{v_{o}^{2}\sin(\theta)\cos(\theta)}{16}\] Follow the same steps as above \[\rightarrow \sin(2\theta) = 0.888\] \[2\theta = 62.73\] \[\theta = 31.365\]